# From Tower to Spire: Adding the Speech Modality to a Text-Only LLM

### v1

@@ Appendix B CTC-based ASR model
We train a CTC-based ASR model using the HuggingFace Transformers library Wolf (2019), leveraging the ASR data from the IT stage (CommonVoice, Table 2) as training data. Our ASR model is made of the HuBERT-Large141414https://huggingface.co/facebook/hubert-large-ll60k speech representation model, followed by three hidden layers and a vocabulary projection layer. We train for 50 epochs with a dropout of 0.3 and a learning rate of 1e-4 with a warm-up ratio of 0.15. The best checkpoint is selected using CER scores. This was obtained at step 220K (at epoch 12.8).
Appendix C Translation-related Tasks
Tower was evaluated on translation-related tasks in addition to MT. We follow their evaluation setup and use the Tower-eval suite151515https://github.com/deep-spin/tower-eval Alves et al. (2024) to evaluate Spire models on APE and NER. For a detailed description of the tasks, we refer the readers to Alves et al. (2024). Briefly, APE measures final translation quality on WMT23 after post-editing with NLLB-3.3B Team et al. (2022), and NER measures entity recognition on MultiCoNER 2023 Fetahu et al. (2023) test. We report COMET-22 for APE and sequence F1 score for NER. We evaluate APE for en{de, ru, zh} and NER for {de, en, es, fr, it, pt, zh}.
Table 8 reports our results. We achieve comparable scores to Towerinstruct across both tasks and all language directions considered. Notably, we outperform TowerFull in all settings, which may be suggestive of the benefit of including text data in the CPT stage.
| APE | NER | ||
| enxx | xxen | Multilingual | |
| TowerInstruct-7b | 83.08 | 80.29 | 71.56 |
| TowerFull | 82.65 | 79.90 | 65.07 |
| SpireFull | 83.13 | 80.08 | 67.10 |

### v3

@@ A.1 Speech Data Preprocessing
A.2 Prompt Format
Table 11 show the prompts used during both training stages.
| ASR (CPT) | |||
| MT (CPT) | |||
| ASR (IT) | |||
| Direct ST (IT) | |||
| Multi-turn ST (IT) | |||
@@ Appendix B CTC-based ASR model
We train a CTC-based ASR model using the HuggingFace transformers library Wolf et al. (2019), leveraging the ASR data from the IT stage (CV, Table 1) as training data after whisper normalization. Our ASR model is made of the HuBERT-Large171717https://huggingface.co/facebook/hubert-large-ll60k speech representation model, followed by three hidden layers and a vocabulary projection layer. We train for 50 epochs with a dropout of 0.3 and a learning rate of 1e-4 with a warm-up ratio of 0.15. We perform step-based best checkpoint selection based on CER scores. Our best checkpoint was obtained at step 220K (at epoch 12.8).
Appendix C ST results
Table 12 report results of ST on FLEURS across baseline models and SpireFull. We report COMET-22. We observe the same trend in scores as reported by spBLEU where in SpireFull obtains the best self-cascaded performance while beating Qwen2-Audio and SALMONN on direct ST across most language pairs. SeamlessM4T obtains the overall best performance in direct ST.
| de | es | fr | it | ko | nl | pt | ru | zh | |||
| Self-Cascade | |||||||||||
| SeamlessM4T | 72.69 | 76.97 | 78.06 | 76.03 | 75.33 | 72.58 | 78.25 | 79.38 | 69.76 | 74.91 | 75.45 |
[…]
| Direct | |||||||||||
| SeamlessM4T | 84.79 | 83.20 | 85.32 | 85.03 | 85.17 | 85.17 | 86.75 | 86.31 | 79.90 | 84.31 | 84.63 |
| SALMONN | 77.41 | 77.99 | 79.95 | 74.47 | 61.07 | 77.18 | 80.94 | 53.05 | 81.63 | 78.51 | 73.74 |
| Qwen2-Audio | 79.82 | 80.43 | 79.44 | 81.28 | 69.33 | 78.75 | 83.41 | 77.90 | 80.71 | 80.55 | 79.01 |
[…]
Appendix D Ablation results
Table 13 reports results from all remaining evaluation datasets across ASR, MT, and ST. We report the same metrics as in Section 4. Here as well, we note that in MT, the inclusion of speech data did not degrade text-only performance (SpireFull vs. TowerFull). Similarly, the inclusion of task-specific text data also did not harm performance on ASR (SpireNoBlocks vs. SpireFull). Lastly, SpireFull has the best performing direct ST system, further showing that individual task competencies (in MT and ASR) do not contribute directly to a compositional task (ST) but rather the inclusion of task-specific data leads to the highest gains (SpireNoPseudo vs SpireFull).
| ASR | MT | ST | |||||||
| WER | C22 | spB | C22 | spB | C22 | spB | |||
| LS Other | Fleurs | VoxPopuli | enxx | xxen | enxx | ||||
[…]
| TowerFull | 13.8 | 14.3 | 40.7 | 84.19 | 36.9 | 82.25 | 35.6 | 71.52 | 20.1 |
| SpireNoBlocks | 7.4 | 10.4 | 15.8 | 73.12 | 26.9 | 74.78 | 25.1 | 74.02 | 23.2 |
| SpireNoPseudo | 7.3 | 11.1 | 14.3 | 83.93 | 36.9 | 82.50 | 35.9 | 59.88 | 6.8 |